# VecTrans: LLM Transformation Framework for Better Auto-vectorization on High-performance CPU

### v1

@@ Abstract.
Large language models (LLMs) have demonstrated great capabilities in code generation, yet their effective application in compiler optimizations remains an open challenge due to issues such as hallucinations and a lack of domain-specific reasoning. Vectorization, a crucial optimization for enhancing code performance, often fails because of the compiler’s inability to recognize complex code patterns, which commonly require extensive empirical expertise. LLMs, with their ability to capture intricate patterns, thus providing a promising solution to this challenge. This paper presents VecTrans, a novel framework that leverages LLMs to enhance compiler-based code vectorization. VecTrans first employs compiler analysis to identify potentially vectorizable code regions. It then utilizes an LLM to refactor these regions into patterns that are more amenable to the compiler’s auto-vectorization. To ensure semantic correctness, VecTrans further integrates a hybrid validation mechanism at the intermediate representation (IR) level. With the above efforts, VecTrans combines the adaptability of LLMs with the precision of compiler vectorization, thereby effectively opening up the vectorization opportunities. Experimental results show that among all 50 TSVC functions unvectorizable by Clang, GCC, and BiShengCompiler, VecTrans successfully vectorizes 23 cases (46%) and achieves an average speedup of 2.02x, greatly surpassing state-of-the-art performance.
1. Introduction
@@
Vectorization is an critical compiler optimization that transforms scalar operations into vectorized forms to exploit the parallel processing capabilities of SIMD (Single Instruction, Multiple Data) architectures. Vectorization dramatically improves computational throughput by simultaneously processing multiple data elements as a single vector operation, making it indispensable for performance-critical domains such as scientific computing, machine learning and multimedia processing. The theoretical foundation of vectorization dates back to Flynn’s taxonomy of parallel architectures(flynn2005very, ), but its practical implementation has evolved alongside advancements in processor design, particularly with the proliferation of SIMD instruction sets such as Intel AVX(IntelAVX2, ), ARM SVE(stephens2017arm, ), and RISC-V V Extension(cui2023risc, ). Despite decades of researching and engineering investments, automatic vectorization still faces serious challenges, including complex data dependencies and diverse code patterns.
@@
Currently, several strategies are available to address non-vectorized loops, including manually modifying the source code, using intrinsic instructions, and applying targeted vectorization optimizations. While these methods can be effective, they each come with their own set of drawbacks. Manual source code modifications (e.g., rewriting loops into SIMD-friendly structures) preserve readability through standard C/C++ implementations, enabling cross-platform compatibility via compiler portability. However, such hand-tuning demands significant developer effort and risks introducing subtle errors like incorrect dependency handling. Compiler intrinsics, e.g., Intel’s _mm512_load_ps, provide precise hardware control, but sacrifice portability by binding code to specific instruction sets (e.g., AVX-512 intrinsics fail on ARM platforms) and obfuscating algorithmic intent. Targeted optimizations automate some vectorization tasks but struggle with non-affine loops or dynamic control flow.
@@
Even compiler-driven approaches confront fundamental tensions: while vanilla C/C++ code retains platform independence, reliance on compiler auto-vectorization leads to inconsistent results, where LLVM(lattner2004llvm, )’s Loop Vectorizer succeeds in some cases, and GCC’s Vectorizer performs better in others. The core challenge lies in conflicting priorities: maintaining portable source code requires abstracting hardware details, whereas maximizing SIMD utilization demands deep architectural awareness (e.g., register widths, masking support). This dichotomy forces developers into a lose-lose choice: either accept suboptimal performance with portable scalar code or lock into vendor-specific intrinsics for marginal gains. Moreover, the inherent limitations of heuristic-driven nature of modern compiler vectorization often fails to fully exploit the available SIMD parallelism, leaving significant performance potential untapped.
@@
These limitations collectively highlight the need for a paradigm that transcends the manual-vs-automatic divide—one that preserves code portability while achieving optimizations close to hand-tuned intrinsics. Our work addresses this challenge by leveraging LLMs to infer architectural constraints from code semantics, enabling automated vectorization that adapts to both program logic and hardware capabilities without compromising readability or portability. Pretrained on vast code corpora, LLMs exhibit emergent capabilities in understanding complex program semantics and generating context-aware transformations. Unlike rule-based compilers, LLMs can reason out implicit parallelism opportunities in loops with irregular structures, such as those containing conditional branches or indirect memory accesses. Our approach employs a multi-feedback architecture that simultaneously analyzes data dependencies, predicts vectorization viability, and generates optimized loop variants that can maximize the potential of LLM-driven code transformation.
@@
However, naively applying LLMs to code optimization presents new challenges. Firstly, LLM is a category of probabilistic models proficient in solving some combinatorial explosion problems. Nevertheless, it suffers from serious hallucinations especially in the tasks related to strong reasoning like mathematical or code generation problems. These unresolved questions make its capability unreliable(zhou2024larger, ; mirzadeh2024gsm, ) if utilized in our studied auto-vectorization optimization. To make LLM’s ability more robust and deterministic, a solid framework usually called AI agent is essential to enhance the interaction of LLM with other robust and precise tools like traditional compilers. Secondly, how to generate both syntactically and semantically correct code is usually a fundamental requirement for auto-vectorization in traditional compilers, but the situation is not suited to naive LLMs. To overcome these challenges, we designed a verification-aware multi-feedback iterative refinement (VAMF-IR) framework, also able to viewed as an AI Agent, for auto-vectorization optimization. The iterative refinement method is proposed by Wilkinson(wilkinson1965rounding, ) in linear system solving problems. Simultaneously, it is firstly introduced in LLM by Madaan(madaan2023self, ) to remarkably enhance the capability of naive LLMs using self-refine mechanism Correspondingly, VAMF-IR is divided into three phases: 1) generate the candidate transformed source code using LLM model; 2) verify semantic equivalence between the original and transformed source codes utilizing unit test and, successively, formal verification with a symbolic execution engine; 3) invalid candidates trigger model retraining with counterexamples, progressively improving vectorization accuracy. This feedback loop enables the system to surpass the coverage limitations of static analysis while maintaining correctness guarantees, marking a critical advancement over existing compiler-based approaches.
@@
  • Cross-platform verification: Our method ensures correctness across platforms by performing formal verification at the intermediate representation (IR) level, eliminating the need for platform-specific validation.
  • Effective results: The evaluations show that VecTrans can effectively extend the vectorization cases on the representative TSVC_2, and further achieves significant performance improvements.
void s1113(int iterations, int LEN_1D, float* a, float* b)
for (int nl = 0; nl < 2*iterations; nl++) {
@@
(a) Original code
void s1113_opt(int iterations, int LEN_1D, float* a, float* b) {
for (int nl = 0; nl < 2*iterations; nl++) {
@@
(b) Optimized code
Figure 1. An example from TSVC_2. LLVM fails to vectorize the original code in (a) due to dependence analysis, restructuring the code in (b) allows the compiler to successfully perform auto-vectorization.
2. Background and Motivation
2.1. Background
@@
Modern compilers generally employ two primary vectorization strategies, namely loop vectorization and Superword Level Parallelism (SLP) vectorization(larsen2000exploiting, ). Loop vectorization, the more established approach, transforms entire loop iterations into vector operations by exploiting cross-iteration parallelism. However, this method requires an extra verification given that loop-carried dependencies are absent. In contrast, SLP vectorization operates at the basic block level, identifying and grouping isomorphic scalar operations within loop bodies or straight-line code into vector instructions (porpodas18lslp, ; porpodas18vwslp, ; porpodas19snslp, ). Both vectorization approaches face significant limitations when dealing with real-world complex codes. Loop vectorization inherently fails to handle loop-carried dependencies situations where iteration i+1 depends on results from iteration i (e.g., recurrence relations like prefix sums). SLP vectorization, while less constrained by cross-iteration dependencies, struggles with non-isomorphic operation patterns.
@@
Several studies have explored the application of LLMs in code optimization. LLM-Vectorizer(10.1145/3696443.3708929, ) demonstrates notable performance by transforming source code into vectorized implementations through intrinsic instructions. However, it exhibits three limitations: 1) inadequate handling of loop-carried dependencies that hinder effective parallelization, 2) generation of platform-specific constructs that compromise portability across architectures, and 3) production of low-readability code lacking maintainability. In parallel, Self-Refine(madaan2023self, ) employs iterative LLM-based refinement to progressively enhance code quality. Nevertheless, this approach suffers from insufficient integration with compiler semantics, resulting in suboptimal optimizations due to the absence of compiler-guided metrics and target-specific optimization objectives. Both paradigms reveal fundamental challenges in integrating LLMs with compiler optimization frameworks.
Figure 2. Overview of VecTrans. When a user submits source code requesting vectorization, if the target function is identified as non-vectorizable, the framework initiates an iterative optimization process. Each iteration comprises two phases: feedback collection and refinement. During the feedback phase, the framework aggregates three categories of diagnostic information: self-evaluation feedback from large language models, static analysis feedback from compilers, and runtime validation feedback. The refinement phase then leverages these diagnostic insights to perform semantics-preserving transformations aimed at enhancing vectorization potential. The framework evaluates the optimized function’s vectorizability, proceeding to subsequent iterations if residual vectorization barriers persist with remaining optimization potential. This iteration process terminates either upon successful vectorization achievement or upon reaching optimization convergence.
2.2. Motivation
@@
While compilers have made significant progress in auto-vectorization, they still face fundamental limitations in recognizing and transforming certain code patterns. Even for recently pioneering work related to LLM-Vectorizer(10.1145/3696443.3708929, ) which tries to integrate LLM, compiler and verification into an AI agent framework, it also loses many optimization opportunities possibly due to the non-optimal interaction manner among different agent components.
@@
To better illuminate the motivation of our work, a representative example is illustrated in Figure 1. The original source code as given in Figure 1(a) cannot be vectorized with traditional compilers even on O3 mode for two reasons: (1) when i=LEN_1D/2, a[LEN_1D/2] is updated, which results in a read-after-write (RAW) dependency in subsequent iterations, and (2) simultaneous reading and writing an array a within the loop makes vectorization unsafe, causing the compiler to reject the transformation. Upon further analysis, it is found that a[LEN_1D/2] is a variable and is updated only once, making it loop-invariant when the loop is split into two separate sections. Unfortunately, LLM-vectorizer(10.1145/3696443.3708929, ) also fails to generate the code able to be vectorized even after several iterations. This is mostly because the generated SIMD intrinsic code cannot pass the verification and also cannot be corrected. It indicates that we need to seek for more opportunities on source code level rather than low-level code like intrinsic or assembly, because general LLM are usually well trained on source code datasets.
Motivated by the above experimental and analyzed results, we believe that the interaction manner among LLM, compilers and other tools makes a difference to overall optimization results for auto-vectorization. Corresponding, in our current work, VecTrans framework is developed to resolve or massively improve the interaction behaviors of different components in source code levels. It integrates different advantages for different components including LLM, compilers and verification tools into a multi-feedback framework - VAMF-IR. It means that LLM is responsible for finding creative opportunities because of its high-level correlation analysis strength, while compilers and verification tools are usually conservatively designed, they are more appropriate to provide some accurate feedback information. Besides, they are more suitable to compile source code precisely on backend SIMD instructions than LLM. The further description about VecTrans are presented in the next section.
3. Design
@@
VecTrans is an iterative refinement LLM optimization framework that uses feedback-refined iteration structures(madaan2023self, ). The feedback-refined iteration is a cyclical optimization process comprising two phases: the feedback phase and the refine phase. During the feedback phase, information relevant to the optimized function is gathered, referred to as feedback information. This feedback is utilized to assess the correctness of the optimized function and to identify the potential for vectorization. In the refine phase, this feedback is applied to update and enhance the optimized function. The feedback information is derived from three sources, as illustrated in Figure 2:
@@
  • Source 2: Compiler Feedback. Feedback information from the compiler, including compilation information and vectorization information. This information is used to help the LLM determine whether the currently generated optimized code can be compiled and vectorized successfully.
@@
  • Source 3: Verification Feedback. The feedback information is used to check whether the generated optimized code can pass the unit tests and formal verification. This helps the large model determine the semantic correctness of the optimized code.
3.1. Source 1: Self-Feedback
The self-feedback information is derived from the analytical capabilities of the LLM. Specifically, the model is leveraged to assess various aspects of the currently optimized code, including lexical correctness, syntactic accuracy, semantic integrity, and the effectiveness of vectorization. By conducting a comprehensive evaluation, the LLM generates valuable insights that inform the next steps in the optimization process. The analysis results produced by the model subsequently guide the refinement phase, where the code undergoes further enhancement based on the feedback provided.

### v2

# VecTrans: Enhancing Compiler Auto-Vectorization through LLM-Assisted Code Transformations

@@ Abstract
Auto-vectorization is a fundamental optimization for modern compilers to exploit SIMD parallelism. However, state-of-the-art approaches still struggle to handle intricate code patterns, often requiring manual hints or domain-specific expertise. Large language models (LLMs), with their ability to capture intricate patterns, provide a promising solution, yet their effective application in compiler optimizations remains an open challenge due to issues such as hallucinations and a lack of domain-specific reasoning. In this paper, we present VecTrans, a novel framework that leverages LLMs to enhance compiler-based code vectorization. VecTrans first employs compiler analysis to identify potentially vectorizable code regions. It then utilizes an LLM to refactor these regions into patterns that are more amenable to the compiler’s auto-vectorization. To ensure semantic correctness, VecTrans further integrates a hybrid validation mechanism at the intermediate representation (IR) level. With the above efforts, VecTrans combines the adaptability of LLMs with the precision of compiler vectorization, thereby effectively opening up the vectorization opportunities. Experimental results show that among all TSVC functions unvectorizable by GCC, ICC, Clang, and BiSheng Compiler, VecTrans achieves an geomean speedup of 1.77x and successfully vectorizes 24 of 51 test cases. This marks a significant advancement over state-of-the-art approaches while maintaining a cost efficiency of \$0.012 per function optimization for LLM API usage.
Index Terms:
auto-vectorization, large language model, compiler optimization
I Introduction
@@
Vectorization is a critical optimization that transforms scalar operations into vectorized forms to exploit the parallel processing capabilities of SIMD (Single Instruction, Multiple Data) architectures. Vectorization dramatically improves data throughput, making it indispensable for performance-critical domains such as scientific computing, machine learning, and multimedia processing. Consequently, automatic vectorization now represents a standard feature in all mainstream compilers. Despite decades of research and engineering investment, automatic vectorization still faces serious challenges, including complex data dependencies and diverse code patterns [1, 2].
@@
Currently, several strategies are available to address code vectorization, especially non-vectorized loops, including manual source code modifications, intrinsic usage, and targeted vectorization optimizations. These methods, while functional, are individually constrained by specific drawbacks. Manual source code modifications (e.g., rewriting loops into SIMD-friendly structures) preserves readability through standard C/C++ implementations, enabling cross-platform compatibility via compiler portability. However, manually optimizing programs is not only labourious, but also error-prone. Intrinsics, e.g., Intel’s _mm512_load_ps, provide precise hardware control, but sacrifice portability by binding code to specific instruction sets (e.g., AVX-512 intrinsics are incompatible with ARM platforms) and obfuscating algorithmic intent. Targeted optimizations automate some vectorization tasks but struggle with non-affine loops or dynamic control flow.
@@
Even compiler-driven approaches confront fundamental tensions: while vanilla C/C++ code retains platform independence, reliance on compiler (e.g., GCC, ICC, Clang[3]) auto-vectorization may lead to inconsistent results. The core challenge lies in conflicting priorities: maintaining portable source code requires abstracting hardware details, whereas maximizing SIMD utilization demands deep architectural awareness (e.g., register widths, masking support). This dichotomy forces developers into a choice: either accept suboptimal performance with portable scalar code or lock into vendor-specific intrinsics for marginal gains. Moreover, the inherent limitations of the heuristic-driven nature of modern compiler vectorization often fail to fully exploit the available SIMD parallelism, leaving significant performance potential untapped.
@@
These limitations collectively highlight the need for a disruptive paradigm that transcends the manual-vs-automatic divide—one that preserves code portability while achieving optimizations close to hand-tuned intrinsics. LLMs, pretrained on vast code corpora, exhibit emergent capabilities in understanding complex program semantics and generating context-aware transformations. Unlike rule-based compilers, LLMs can reason out implicit parallelism opportunities in loops with irregular structures, such as those containing conditional branches or indirect memory accesses.
However, applying LLMs to code optimization presents new challenges. Firstly, LLM is a category of probabilistic models proficient in solving some combinatorial explosion problems. Nevertheless, it suffers from serious hallucinations, especially in tasks related to strong reasoning like mathematical or code generation problems. This unresolved question makes its capability unreliable [4, 5] when utilized in previous targeted auto-vectorization. To make LLM’s ability more robust and deterministic, a framework named AI agent is essential to enhance the interaction of LLM with other robust and precise tools like traditional compilers. Secondly, generating both syntactically and semantically correct code is usually a fundamental requirement for auto-vectorization in traditional compilers, but this is not guaranteed in naive LLMs.
@@
To overcome these challenges, we designed a verification-aware multi-feedback iterative refinement framework, which can be considered as an AI agent, for auto-vectorization optimization. The iterative refinement method is proposed by Wilkinson[6] in linear system solving problems. It is firstly introduced in LLM by Madaan[7] to remarkably enhance the capability of naive LLMs using self-refine mechanism. Correspondingly, the framework is divided into three phases: (1) generate the candidate transformed source code using LLM, with the requirement that the generated code is more amenable to compiler auto-vectorization; (2) verify semantic equivalence between the original and transformed source codes utilizing unit test and, successively, conduct the formal verification with a symbolic execution engine; (3) invalid candidates trigger model retraining with counterexamples, progressively improving vectorization accuracy. This feedback loop enables the system to surpass the coverage limitations of static analysis while maintaining correctness guarantees, marking a critical advancement over existing compiler-based approaches. Additionally, the objective of this task is to generate code that is more conducive to compiler auto-vectorization by performing source-level transformations, while ensuring the optimized code maintains readability, portability, and verifiability.
@@
  • Effective results and cross-platform verification: The evaluations show that VecTrans effectively extends the vectorization cases, and further achieves significant performance improvements. Meanwhile, correctness across platforms is ensured by performing formal verification at the IR level, eliminating the need for platform-specific validation.
void s1113(...) {
for (int nl = 0; nl < 2 * iters; nl++)
@@
(a) Original non-vectorizable code.
void s1113_opt(...) {
for (int nl = 0; nl < 2 * iters; nl++){
@@
(b) Transformed vectorizable code generated by VecTrans.
Figure 1: Example from TSVC_2. The original loop cannot be vectorized due to a cross-iteration dependency on a[LEN_1D/2]. VecTrans transforms the loop by splitting it at the dependent index and introducing temporary variables, enabling safe vectorization.
II Background and Motivation
II-A Background
The theoretical foundation of SIMD and vectorization dates back to Flynn’s taxonomy of parallel architectures[8], and its practical implementation has evolved alongside advancements in processor design, particularly with the proliferation of SIMD instruction sets such as Intel AVX[9], ARM SVE [10], and RISC-V V Extension[11]. Vectorization plays a pivotal role in performance-critical domains by dramatically improving instruction-level parallelism. Consequently, automatic vectorization has become an indispensable component of contemporary compiler toolchains.
@@ II-A Background
Two primary vectorization strategies are employed by mainstream compiler toolchains, namely loop vectorization and Superword Level Parallelism (SLP) vectorization[12]. Loop vectorization, the more established approach, transforms entire loop iterations into vector operations by exploiting cross-iteration parallelism. However, this method requires an extra verification given that loop-carried dependencies are absent. In contrast, SLP vectorization operates at the basic block level, identifying and grouping isomorphic scalar operations within loop bodies or straight-line code into vector instructions [13, 14, 15]. Both approaches face significant limitations when dealing with real-world complex codes. Loop vectorization inherently fails to handle loop-carried dependencies, situations where iteration i+1 depends on results from iteration i (e.g., recurrence relations like prefix sums). SLP vectorization, while less constrained by cross-iteration dependencies, struggles with non-isomorphic operation patterns.
@@
The emergence of large language models (LLMs) and their ability to capture complex patterns have spurred research into LLM-based code optimization techniques. LLM-Vectorizer[16] demonstrates notable performance by transforming source code into vectorized implementations through intrinsic instructions. However, three limitations are exhibited: (1) inadequate handling of loop-carried dependencies that hinder effective parallelization, (2) generation of platform-specific constructs that compromise portability across architectures, and (3) production of low-readability code lacking maintainability. In parallel, Self-Refine[7] employs iterative LLM-based refinement to progressively enhance code quality. Nevertheless, this approach suffers from insufficient integration with compiler semantics, resulting in suboptimal optimizations due to the absence of compiler-guided metrics and target-specific optimization objectives. Both paradigms reveal fundamental challenges in integrating LLMs with compiler optimization frameworks. AlphaEvolve[17] also adopts an iterative approach to generate answers while possessing verification capabilities. Its main difference from Self-Refine lies in not dividing the iteration process into distinct feedback and refine phases. Additionally, AlphaEvolve similarly lacks the utilization of tools such as compilers.
Figure 2: The workflow of VecTrans: ❶ Dual-Phase Iteration - Alternating feedback and refinement; ❷ Tri-Source Diagnostics - LLM self-evaluation, compiler-derived static analysis, and runtime validation; ❸ Guided Code Transformation - Employing transformations in the refinement phase informed by diagnostics to improve vectorization compatibility; ❹ Adaptive Termination - Dynamic termination when reaching either successful vectorization or optimization thresholds.
II-B Opportunities of LLM-assisted code transformation
@@
While compilers have made significant progress in auto-vectorization, they still face fundamental limitations in recognizing and transforming certain code patterns. As shown in Table I, 50 out of 149 functions in the TSVC benchmark resist compiler auto-vectorization. Even for recently pioneering work related to LLM-Vectorizer[16] –which tries to integrate LLM, compiler, and verification into an unified AI agent framework, it also miss substantial optimization opportunities, due to the non-optimal interaction manner among different agent components.
@@
To illuminate the missed opportunity of auto-vectorization by code transformation, we present an example in Figure 1. The original source code as given in Figure 1(a) cannot be vectorized with traditional compilers even on -O3 mode for two reasons: (1) when i=LEN_1D/2, a[LEN_1D/2] is updated, which results in a read-after-write (RAW) dependency in subsequent iterations, and (2) simultaneous reading and writing an array a within the loop makes vectorization unsafe, causing the compiler to reject the transformation. Upon further analysis, it is found that a[LEN_1D/2] is a variable and is updated only once, making it loop-invariant when the loop is split into two separate sections. Our evaluation reveals that LLM-Vectorizer[16] fails to produce vectorizable code across multiple iterations. This stems from two key limitations: (1) generated SIMD intrinsics consistently fail verification, and (2) the system lacks effective correction mechanisms. These findings suggest that LLM-based vectorization may benefit more from source-level transformations than low-level code generation, as LLMs appear more adept at handling high-level code patterns derived from their training data.
We summarize three key challenges in LLM-based auto-vectorization: (1) Suboptimal Component Interaction: Naive integration of LLMs with compiler toolchains leads to verification failures (Missing compiler feedback roughly halves performance) and missed optimization opportunities due to incoherent feedback loops. (2) Semantic-Preservation Dilemma: LLMs’ high-level pattern matching often violates compiler-required semantics (e.g., aliasing rules), while strict verification overly constrains optimization potential. (3) Large language models manifest erratic convergence behavior during generative refinement iterations (e.g., oscillating between excessive vectorization attempts and conservative rollbacks), whereas traditional compiler heuristics lack the adaptive mechanisms needed to orchestrate their co-evolution toward convergent optimizations.
III Design
@@
In this section, we propose VecTrans, an iterative refinement LLM optimization framework that uses feedback-refined iteration structures[7]. The feedback-refined iteration is a cyclical optimization process comprising two phases: the feedback phase and the refine phase.
During the feedback phase, information relevant to the optimized function is gathered, referred to as feedback information. This feedback is utilized to assess the correctness of the optimized function and to identify the potential for vectorization. The feedback information is derived from three sources, as illustrated in Figure 2:
@@
  • Source 2: Compiler Feedback. Feedback information from the compiler, including compilation information and vectorization information. This information is used to determine whether the currently generated optimized code can be compiled and vectorized successfully.
@@
  • Source 3: Verification Feedback. The feedback information is used to check whether the generated optimized code can pass the unit tests and formal verification. This helps the large language model determine the semantic correctness of the optimized code.
Following the feedback phase, a structured prompting strategy is applied during the refine phase to synthesize multi-source feedback into actionable optimizations. Specifically, the three feedback streams are systematically integrated with a vectorization-focused prompt template to guide the LLM’s code regeneration.
III-A Self-Feedback
The self-feedback mechanism harnesses the LLM’s analytical capabilities to evaluate optimized code along four dimensions: lexical correctness, syntactic validity, semantic equivalence, and vectorization potential. Through structured prompting, the LLM conducts a staged analysis process: (1) Lexical and Syntactic Validation: detecting tokenization and grammar errors; (2) Semantic Equivalence Checking: simulating unit test outcomes to predict runtime behavior without actual execution; (3) Vectorization Feasibility Assessment: analyzing loop structures to determine whether full vectorization is achievable; and (4) Optimization Opportunity Diagnosis: identifying targeted transformations for semantically correct code. This comprehensive evaluation yields actionable insights that directly inform iterative optimization decisions.